# competing interactions; square well; square shoulder; discontinuous structural crossover transitions; Fisher–Widom line; Rational-Function Approximation

### v1

@@ A.1 Absence of real poles if $\varphi(r)\geq 0$
We first consider a generic potential $\varphi(r)$ that goes to $\infty$ if $r<\sigma$ and vanishes if $r>b>\sigma$. If a real pole $s=-\kappa<0$ of $G(s)$ exists, then
@@
In summary, if $\varphi(r)\geq 0$ for all $r$, then no real poles exist. Otherwise, real poles $s=-\kappa$ may exist but, then, $\kappa>\beta p$.
A.2 Poles in the high-temperature limit
In the limit $\beta\to 0$, the system simplifies to a HR fluid with diameter $\sigma=1$. In this regime, Equations (9) and (10) reduce to
@@ A.2 Poles in the high-temperature limit
In the close-packing limit $\rho^{*}\to 1$, it is easy to obtain
@@
On the other hand, in the opposite low-density limit ($\rho^{*}\to 0$) one has $\omega_{\text{HR}}\approx\pi$ and $\zeta_{\text{HR}}e^{-\zeta_{\text{HR}}}=\rho^{*}$. The latter belongs to the class of Lambert equations $z=we^{w}$ with $z=-\rho^{*}$ and $w=-\zeta_{\text{HR}}$. The solution is then
@@
where $W_{-1}(z)$ is the lower branch of the Lambert function Corless et al. (1996). The functions $\zeta_{\text{HR}}(\rho^{*})$ and $\omega_{\text{HR}}(\rho^{*})$ are plotted in Figure 8
@@
A.3 Poles in the low-temperature limit with $\epsilon_{1}>0$
@@ A.3.1 Case $0<\rho^{*}<\lambda_{2}^{-1}$
In the low-temperature limit $\beta\to\infty$ with $\epsilon_{1}>0$, one can see from Equations (7)–(10) that the system becomes equivalent to an HR fluid with diameter $\lambda_{2}$, provided that $\rho^{*}<\lambda_{2}^{-1}$. Therefore, in that limit one has
@@ A.3.2 Case $\lambda_{2}^{-1}<\rho^{*}<1$
In the limit $\beta\to\infty$ with finite $p$ one has $e^{-\beta p}\gg e^{-\beta p\lambda_{1}}\gg e^{-\beta p\lambda_{2}}$, so that Equation (9) yields
@@
Suppose first that $0<\epsilon_{1}/\epsilon_{2}\leq(\lambda_{2}-1)/(\lambda_{2}-\lambda_{1})$. In that case, the inequalities in (28) and (29) imply $p<\epsilon_{1}/(\lambda_{2}-1)$ and $p>\epsilon_{1}/(\lambda_{2}-1)$, respectively, while the inequality in (30) is impossible. Thus, in the low-temperature limit, the density changes from $\rho^{*}=\lambda_{2}^{-1}$ to $\rho^{*}=1$ as pressure crosses the value $p=\epsilon_{1}/(\lambda_{2}-1)$. To analyze this change in detail, let us introduce the scaled variable $\mu$ by
@@
In turn, from Equations (10) we have
@@
For simplicity, let us assume that $\lambda_{2}$ is a rational number $\lambda_{2}=m_{2}/n_{2}$. The analytical solution to Equations (33) in the limit $\beta\epsilon_{1}\to\infty$ is displayed in the first row of Table 1.
Table 1: Asymptotic expressions for $p$, $\zeta$, and $\omega$ in the low-temperature limit for 1D fluids with $\epsilon_{j}>0$, assuming $\lambda_{1}=m_{1}/n_{1}$, $\lambda_{2}=m_{2}/n_{2}$, and $\lambda_{2}/\lambda_{1}=m_{21}/n_{21}$ are rational numbers.
@@
In contrast, for a potential where $\epsilon_{1}/\epsilon_{2}>(\lambda_{2}-1)/(\lambda_{2}-\lambda_{1})$, the situation becomes more intricate. In this case, the inequalities in Equations (28), (29), and (30) imply the following conditions: $p<\epsilon_{2}/(\lambda_{2}-\lambda_{1})$, $p>(\epsilon_{1}-\epsilon_{2})/(\lambda_{1}-1)$, and $\epsilon_{2}/(\lambda_{2}-\lambda_{1})<p<(\epsilon_{1}-\epsilon_{2})/(\lambda_{1}-1)$, respectively. This indicates that, in the low-temperature limit, the density changes from $\rho^{*}=\lambda_{2}^{-1}$ to $\rho^{*}=\lambda_{1}^{-1}$ as the pressure crosses $p=\epsilon_{2}/(\lambda_{2}-\lambda_{1})$, and from $\rho^{*}=\lambda_{1}^{-1}$ to $\rho^{*}=1$ as the pressure crosses $p=(\epsilon_{1}-\epsilon_{2})/(\lambda_{1}-1)$. An analysis similar to that conducted in Equations (31)–(33) leads to the expressions shown in the second and third rows of Table 1. An oscillation discontinuity at $\rho=\lambda_{1}^{-1}$ only arises if $n_{21}\neq m_{1}$, meaning $n_{1}/n_{2}$ must be an integer. This condition excludes cases such as $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$, where $n_{1}=20$, $m_{1}=27$, $n_{2}=10$, $m_{2}=17$, $n_{21}=27$, and $m_{21}=34$.
@@ Appendix B Parameters in Equation (15)
Further, since the cavity function $y(r)\equiv g(r)e^{\beta\varphi(r)}$ must be continuous at $r=\lambda_{1}$ and $r=\lambda_{2}$, the two following conditions should also hold Santos et al. (2012)
@@
This closes the problem of determining the nine parameters in terms of $\eta$, $\lambda_{1}$, $\lambda_{2}$, $\beta\epsilon_{1}$, and $\beta\epsilon_{2}$. In fact, Equations (34) allow us to express $S_{1}$, $S_{2}$, $S_{3}$, and $B_{0}$ as linear combinations of $B_{1}$ and $B_{2}$, so that in the end one only has to solve (numerically) the two coupled transcendental equations (36).

### v2

@@ A.1 Absence of Real Poles If $\varphi(r)\geq 0$
We first consider a generic potential, $\varphi(r)$, that goes to $\infty$ if $r<\sigma$ and vanishes if $r>b>\sigma$. If a real pole, $s=-\kappa<0$, of $G(s)$ exists, then
@@
In summary, if $\varphi(r)\geq 0$ for all $r$, then no real poles exist. Otherwise, the real poles $s=-\kappa$ may exist and, then, $\kappa>\beta p$.
A.2 Poles for the High-Temperature Limit
For the limit $\beta\to 0$, the system simplifies to an HR fluid with the diameter $\sigma=1$. In this regime, Equations (9) and (10) reduce to
@@ A.2 Poles for the High-Temperature Limit
For the close-packing limit $\rho^{*}\to 1$, it is easy to obtain
@@
On the other hand, for the opposite low-density limit ($\rho^{*}\to 0$), one has $\omega_{\text{HR}}\approx\pi$ and $\zeta_{\text{HR}}e^{-\zeta_{\text{HR}}}=\rho^{*}$. The latter belongs to the class of Lambert equations $z=we^{w}$ with $z=-\rho^{*}$ and $w=-\zeta_{\text{HR}}$. The solution is then
@@
where $W_{-1}(z)$ is the lower branch of the Lambert function Corless et al. (1996). The functions $\zeta_{\text{HR}}(\rho^{*})$ and $\omega_{\text{HR}}(\rho^{*})$ are plotted in Figure 9.
@@
A.3 Poles for the Low-Temperature Limit with $\epsilon_{1}>0$
@@ A.3.1 Case $0<\rho^{*}<\lambda_{2}^{-1}$
For the low-temperature limit $\beta\to\infty$ with $\epsilon_{1}>0$, one can see from Equations (7)–(10) that the system becomes equivalent to an HR fluid with the diameter $\lambda_{2}$, provided that $\rho^{*}<\lambda_{2}^{-1}$. Therefore, for that limit, one has
@@ A.3.2 Case $\lambda_{2}^{-1}<\rho^{*}<1$
For the limit $\beta\to\infty$ with finite $p$, one has $e^{-\beta p}\gg e^{-\beta p\lambda_{1}}\gg e^{-\beta p\lambda_{2}}$, so Equation (9) yields
@@
Suppose first that $0<\epsilon_{1}/\epsilon_{2}\leq(\lambda_{2}-1)/(\lambda_{2}-\lambda_{1})$. In that case, the inequalities in (28) and (29) imply $p<\epsilon_{1}/(\lambda_{2}-1)$ and $p>\epsilon_{1}/(\lambda_{2}-1)$, respectively, while the inequality in (30) is impossible. Thus, for the low-temperature limit, the density changes from $\rho^{*}=\lambda_{2}^{-1}$ to $\rho^{*}=1$ as the pressure crosses the value $p=\epsilon_{1}/(\lambda_{2}-1)$. To analyze this change in detail, let us introduce the scaled variable $\mu$ by
@@
In turn, from Equation (10), we have
@@
For simplicity, let us assume that $\lambda_{2}$ is a rational number, $\lambda_{2}=m_{2}/n_{2}$. The analytical solution to Equations (33) in the limit $\beta\epsilon_{1}\to\infty$ is displayed in the first row of Table 1.
Table 1: Asymptotic expressions for $p$, $\zeta$, and $\omega$ for the low-temperature limit for 1D fluids with $\epsilon_{j}>0$, assuming $\lambda_{1}=m_{1}/n_{1}$, $\lambda_{2}=m_{2}/n_{2}$, and $\lambda_{2}/\lambda_{1}=m_{21}/n_{21}$ are rational numbers.
@@
In contrast, for a potential where $\epsilon_{1}/\epsilon_{2}>(\lambda_{2}-1)/(\lambda_{2}-\lambda_{1})$, the situation becomes more intricate. In this case, the inequalities in Equations (28)–(30) imply the following conditions: $p<\epsilon_{2}/(\lambda_{2}-\lambda_{1})$; $p>(\epsilon_{1}-\epsilon_{2})/(\lambda_{1}-1)$; and $\epsilon_{2}/(\lambda_{2}-\lambda_{1})<p<(\epsilon_{1}-\epsilon_{2})/(\lambda_{1}-1)$, respectively. This indicates that, for the low-temperature limit, the density changes from $\rho^{*}=\lambda_{2}^{-1}$ to $\rho^{*}=\lambda_{1}^{-1}$ as the pressure crosses $p=\epsilon_{2}/(\lambda_{2}-\lambda_{1})$ and from $\rho^{*}=\lambda_{1}^{-1}$ to $\rho^{*}=1$ as the pressure crosses $p=(\epsilon_{1}-\epsilon_{2})/(\lambda_{1}-1)$. An analysis similar to that conducted in Equations (31)–(33) leads to the expressions shown in the second and third rows of Table 1. An oscillation discontinuity at $\rho=\lambda_{1}^{-1}$ only arises if $n_{21}\neq m_{1}$, meaning that $n_{1}/n_{2}$ must be an integer. This condition excludes cases such as $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$, where $n_{1}=20$, $m_{1}=27$, $n_{2}=10$, $m_{2}=17$, $n_{21}=27$, and $m_{21}=34$.
@@ Appendix B Parameters in Equation (15)
Further, since the cavity function $y(r)\equiv g(r)e^{\beta\varphi(r)}$ must be continuous at $r=\lambda_{1}$ and $r=\lambda_{2}$, the two following conditions should also hold Santos et al. (2012):
@@
This closes the problem of determining the nine parameters in terms of $\eta$, $\lambda_{1}$, $\lambda_{2}$, $\beta\epsilon_{1}$, and $\beta\epsilon_{2}$. In fact, Equation (34) allows us to express $S_{1}$, $S_{2}$, $S_{3}$, and $B_{0}$ as linear combinations of $B_{1}$ and $B_{2}$ so that in the end, one only has to solve (numerically) the coupled transcendental Equation (36).